# Foregrounding Artist Opinions: A Survey Study on Transparency, Ownership, and Fairness in AI Generative Art

### 30 Jan 2024 v2

@@ 2.4. Problems in the pipeline of Generative AI tools
As mentioned, when we consider the impact of an AI model, the impact is not just how the model’s output is used or how the model itself is deployed (as a replacement for human workers, etc.). Each stage within the pipeline of creating a Generative AI model can contribute to the eventual harm done by (and therefore the impact of) the model, even though most stages, with the sometimes exception of the two endpoints (input and output), are relatively inaccessible. In this section, we outline the pipeline of creating a Generative AI model, and we examine one particularly pervasive harm, bias (Zhao et al., 2023; Gallegos et al., 2023), in order to explain how multiple stages within the pipeline can contribute to the model’s impact.
When creating an AI model, training materials must be gathered (often a subset of artifacts from the real world) (Salkowitz, 2022); those materials must be groomed; poor quality, irrelevant, or over-represented artifacts should be removed or otherwise addressed (Luccioni and Viviano, 2021; Zhao et al., 2023). The architecture of the model itself must be assembled. The model must be trained. For complex Generative AI models whose behavior can be unpredictable, there is the additional process of alignment, which is intended to bring the model’s range of possible — or at least likely – behavior in line with whatever the creators of the model deem desirable (Zhao et al., 2023).
@@
Very early in the pipeline, choosing the training material has a significant influence on the model’s eventual behavior. Important datasets used as benchmarks for machine learning have been found to be “overwhelmingly composed of lighter-skinned subjects” (Buolamwini and Gebru, 2018), introducing bias obliquely into the model (the global bias introduced would not be caught by locally vetting each piece of training data for harmful content). However, there is also explicitly objectionable and biased content in the datasets themselves – such as racist hate speech and harmfully sexualized images (Luccioni and Viviano, 2021; Cole, 2023) – even after cleaning procedures.
Midway through the pipeline, in the interior of the trained model, pieces as basic and abstracted as the meaning of common English words have been found to encode gender biases (Bolukbasi et al., 2016). Near the end of the pipeline, the output produced by the model can instantiate bias. For example, the Stable Diffusion model disproportionately generates images of white male executives, doctors, lawyers, judges, and law enforcement, and also disproportionately generates images of black and brown criminals and service workers (Nicoletti and Bass, 2023; Ghosh and Caliskan, 2023). Additionally, some users have found that Generative AI tools readily produced images in which they looked younger, more sexualized, or whiter, even in cases where their prompt did not request or hint at those parameters (Snow, 2022; Ghosh and Caliskan, 2023; Hall, 2023).
These examples are not unique; similar biases are found across current AI models and products – including those of Generative AI – at every level, from the training data they learned from to their internal representations to their outputs (Luccioni and Viviano, 2021; Ghosh and Caliskan, 2023; Cheng et al., 2023; Gallegos et al., 2023; Zhao et al., 2023; Common Crawl, [n. d.]; Cole, 2023).
@@
As a final problem in the pipeline, it is worth noting that these AI tools cannot be created ex nihilo: they are integrated into complex human systems that bring their own forces to bear on the model’s potential impact. There are discrepancies in the demographics of who is wielding the power of these new technologies: who is making them (for Human-Centered Artificial Intelligence, 2023), profiting off them, using them. This can have many knock-on effects, sometimes through mechanisms difficult to foresee (Even-Tov et al., 2023), with respect to the design, use, and role of these tools in society (Jiang et al., 2023; for Human-Centered Artificial Intelligence, 2023). In particular, the companies behind these products often employ a great deal of secrecy as to their training data, architecture, or implementation (or all three) (Ho, 2024; the Midjourney team, 2023). As a result, users must interact with the product as a black box (Jiang et al., 2023; Marcus and Southen, 2024); concerned users can only conjecture as to what kind of training data could produce the observed results based on what is known of the basic architecture of the model, which in turn allows argument over the obscure details (Team, 2024; Masse, 2024). The lack of transparency benefits the companies offering these products (with respect to establishing the impact, specifically the harms, of the model) in giving them a way to sidestep moral and ethical concerns (Marcus and Southen, 2024; Masse, 2024).
2.5. Previous disruption: Napster’s example
Similarities exist between the early years of the Peer-to-Peer music sharing platform Napster and the nascent stage of Generative AI tools. Both technologies were exciting and disruptive, with the potential for financially high stakes (Jiang et al., 2023; for Human-Centered Artificial Intelligence, 2023; Bubeck et al., 2023; Bergen, 2002). Both also forged ahead despite glaring questions of legality, particularly with respect to copyright (Grynbaum and Mac, 2023). In Napster’s wake, legal scholars debated whether contemporaneous copyright law did or should apply to the digital age, and they wondered whether the public’s perception of copyright laws had changed due to Napster’s large user base and the demand for free or reduced-cost music (Ku, 2002; Bergen, 2002). Ultimately, Napster shut down, having underestimated the legal battles it would face (Green, 2002; YellowBrick, 2023).
The argument that Generative AI technologies should be able to be “taught” their downstream abilities (to speak a language, to reason, to create various output) (Masse, 2024; Team, 2024; team, 2023) by much of the (ostensibly (Marcus and Southen, 2024)) publicly available Internet (Salkowitz, 2022) – or that, morality aside, to prevent them would be infeasible – is a more complex version of the argument made by Napster’s founders and early users who insisted that music should (or would) be available to the public for free (Marcus and Southen, 2024).
In the face of such arguments, it is worth noting that openness is not a good in and of itself, and can lead to its own harms, such as violations of contextual integrity or privacy (Nissenbaum, 2004). Furthermore, the forces, often economic, that shape underlying practices like web scraping (Common Crawl, [n. d.]; Salkowitz, 2022) may not be the forces best suited to beneficently shaping technological development (AI, 2023b).
The apparent success of music distribution platforms such as Apple Music and Spotify (Wise, 2023) – and the absence of Napster – suggests that there is significant doubt as to the eventual triumph of Generative AI models over pending legal challenges. As of now, at least, there are no exceptions for AI companies regarding US copyright law (Commission, 2023).
@@ 3. Results
Our anonymous survey of artists is an observational study (N=459) of artists over the age of 18 years (we report participant demographics in Table 1). We see from the responses to our primary research questions (see Table 2) that 80.17% of the survey participants agree that model creators should be required to disclose in detail what art and images they use to train their AI models. We also find that 61.87% of survey participants agree that AI models are a threat to art workers. In contrast, 44.88% of survey participants agree that AI art models are a positive development in the field of art, notably with 27.02% of artists selecting neutral on the Likert scale for this question. Interestingly, we see an overlap between artists who agree that AI art is a positive development and those who also agree that it is a threat to art workers; 22.03% of survey participants agreed that both AI was a positive development and a threat.
We look at the question that asks participants to think about a scenario where an AI art model was used by someone else to produce artwork recognizably in their (the survey participant’s) style and think about who should own the work and its derivatives. Only 26.80% of the participants agreed that the work and its derivatives be considered the property of the AI model creators, while 39.22% of participants agreed that the work and its derivatives be considered the property of the person who used the AI model to generate the artwork. Finally, 41.39% of the participants agreed that the work and its derivatives be considered the property of the artists whose style is being represented by the AI model output. In the following sections, we will analyze which variables are most associated with participants in our survey study who agreed with the primary research questions.
@@
Table 1. Descriptive statistics for our survey participants, N=459.
@@
3.1. RQ1: Do artists see AI models as a threat to art workers?
Our results show that 61.87% of survey participants believe that AI models pose a threat to art workers. This finding is related to survey question no. 15 (seen in Section B). An ordinal logistic regression (OLR) was conducted to investigate the association between the Likert scale answers to RQ1 (dependent variable) and the secondary variables (as shown in Figure 3) in terms of the odds ratio of the OLR.
The results demonstrate a significant association (p-value ¡ 0.05, indicated in Table 4) between agreeing (represents a combination of both the responses ‘Agree’ and ‘Strongly Agree’ in the Likert) that AI models are a threat to art workers and age, as well as whether or not the participants have used AI to create art.
@@ 3.1. RQ1: Do artists see AI models as a threat to art workers?
Figure 3. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ1 (AI models as a threat to art workers) and other survey variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
A forest plot centered on $10^{0}$ that shows the odds ratio and the 95% confidence interval on the x-axis and the variable on the y-axis. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ1 and the survey’s secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Figure 3. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ1 (AI models as a threat to art workers) and other survey variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Secondary Variable Participant Characteristic Primary Variable Likert Question Odds of Agreeing with RQ p-value
Used AI to make art RQ1: AI models are a threat to art workers 1.763 0.005
18-29 years old RQ1: AI models are a threat to art workers 2.329 0.002
Not makers RQ2: AI art models are a positive development in art 13.513 0.047
30-49 years old RQ2: AI art models are a positive development in art 1.535 0.039
Male and non-binary RQ2: AI art models are a positive development in art 1.663 0.007
Male RQ2: AI art models are a positive development in art 1.570 0.012
Not an artists RQ2: AI art models are a positive development in art 1.736 0.019
White RQ2: AI art models are a positive development in art 1.494 0.004
Used AI to make art RQ2: AI art models are a positive development in art 5.178 ¡0.001
Compensation: Would donate their - artwork to train models RQ2: AI art models are a positive development in art 15.444 ¡0.001
18-49 years old and ages above 65 required disclosure from model creators 2.439 0.001
18-29 years old RQ3: required disclosure from model creators 2.874 0.003
30-49 years old RQ3: required disclosure from model creators 1.777 0.028
Familiarity with AI art generally RQ3: required disclosure from model creators 2.620 0.003
Not an illustrator RQ3: required disclosure from model creators 4.822 0.026
Compensation: a portion of any - profit from derivatives RQ3: required to disclose 6.461 0.038
Compensation: portion - of any profit made RQ3: required disclosure from model creators 4.213 0.042
Has not purchased art RQ4_A: AI users owners of AI art made 1.773 0.039
Male RQ4_AA: users owners of AI art made 2.019 ¡0.001
Male and non-binary RQ4_A: I users owners of AI art made 2.197 ¡0.001
Compensation: did not select “not comfortable - with any listed compensation type” RQ4_A: users owners of AI art made 4.764 0.018
Person of color RQ4_B: Artists owners of AI art made 2.147 ¡0.001
Not used AI to make art RQ4_B:Artists owners of AI art made 1.523 0.0255
Person of color RQ4_C: Model creators owners of AI art made 2.491 ¡0.001
Male RQ4_C:Model creators owners of AI art made 2.033 ¡0.001
Male or non-binary RQ4_C: Model creators owners of AI art made 1.949 ¡0.001
Female or male RQ4_C: Model creators owners of AI art made 2.815 0.033
Not have a familiar with AI art generally RQ4_C: Model creators owners of AI art made 3.671 ¡0.001
Tattooists RQ4_C: Model creators owners of AI art made 6.674 0.042
Compensation: None, donate use - of artwork to train models RQ4_C: Model creators owners of AI art made 6.074 0.010
Table 4. Table represents the results of an ordinal logistic regression (OLR) to explore the relationship between our dependent variable (the Likert scale answers to each research question RQ1-4) and our secondary variables. The odds ratio is represented as the likelihood that the participants who selected a particular secondary variable agree or strongly agree with the primary variables research question. For example, for participants who have not used AI to make art, the odds of being more likely to agree that AI models are a threat to art workers is 1.7633 times that of participants who have used AI to make art (p-value=0.005). We consider evidence of a strong association to be a p-value less than 0.05.
3.2. RQ2: Do artists see AI art models as a positive development in the field of art?
We ran the same analysis for our second research question and found that 44.88% of the survey participants agree that AI art models are a positive development in the field of art. Notably, 27.02% of artists selected ’neutral’ on the Likert scale for this question. This analysis pertains to survey question number 14 (see the section on survey questions B). We ran an ordinal logistic regression (OLR) to explore the relationship between the Likert scale answers to RQ2 (the dependent variable) and the secondary variables (as seen in Figure 4, which shows the odds ratio of the OLR).
Our analysis indicates a strong association (p-value ¡ 0.05, as seen in Table 4) between agreeing that AI art models are a positive development in the field of art and various factors such as art practice, age, gender, artistic status, race, whether or not they have used AI to make art and compensation.
@@ 3.2. RQ2: Do artists see AI art models as a positive development in the field of art?
Figure 4. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ2 (AI models as a positive development) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
A forest plot centered on $10^{0}$ that shows the odds ratio and the 95% confidence interval on the x-axis and the secondary variable on the y-axis. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ2 and the survey’s secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Figure 4. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ2 (AI models as a positive development) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
3.3. RQ3: What types of disclosure for the use of artwork to train an AI art model do artists consider fair?
We see from our survey results that 80.17% of the survey participants agree that creators of AI models should disclose the art and images used to train their models in detail. The analysis for RQ3 is related to survey question 13 (seen in Section B), which we used to run an ordinal logistic regression (OLR) to explore the relationship between the Likert scale answers to RQ3 and secondary variables. The results of the OLR, displayed in Figure 5 as the odds ratio, showed a strong association (p-value ¡ 0.05, as seen in Table 4) between agreeing with the disclosure requirement and factors such as art practice, age, familiarity with AI, whether or not they have used AI to create art and compensation preference.
@@
Figure 5. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ3 (Should model creators be required to disclose in detail what art & images they use to train their AI models) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
A forest plot centered on $10^{0}$ that shows the odds ratio and the 95% confidence interval on the x-axis and the secondary variable on the y-axis. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ3 and the survey’s secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Figure 5. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ3 (Should model creators be required to disclose in detail what art & images they use to train their AI models) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
3.4. RQ4: From an artist’s perspective, who should own the derivative artwork created by AI art models?
Finally, we look at the question that asks participants to think about a scenario where an AI art model was used by someone else to produce artwork recognizably in their (the survey participant’s) style and think about who should own the work and its derivatives. The analysis for RQ4 is related to survey questions B No. 8,9 and 10.
RQ4_A: If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh), Should that work and its derivatives be considered the property of the person who used the AI model to generate the artwork? We see that 39.22% of participants agreed that the work and its derivatives be considered the property of the person who used the AI model to generate the artwork. We run an ordinal logistic regression (OLR) to explore the relationship between our dependent variable (the Likert scale answers to RQ4_A) and our secondary variables, as seen in Figure 6 as the odds ratio of the OLR.
@@ 3.4. RQ4: From an artist’s perspective, who should own the derivative artwork created by AI art models?
Figure 6. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4_A (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) Should that work and its derivatives be considered the property of the person who used the AI model to generate the artwork?) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
A forest plot centered on $10^{0}$ that shows the odds ratio and the 95% confidence interval on the x-axis and the secondary variable on the y-axis. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ4_Aand the survey’s secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
Figure 6. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4_A (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) Should that work and its derivatives be considered the property of the person who used the AI model to generate the artwork?) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value ¡ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
We see evidence of a strong association (p-value ¡ 0.05, seen in Table 4) between agreeing that the person who used the AI model to generate the artwork should be the owner, and gender, having purchased art, race, compensation, and having used AI to make art.
RQ4_B: If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh and you are Van Gogh), Should that work and its derivatives be considered yours?
Our survey shows that 41.39% of the participants agreed that the work and its derivatives should be considered their property (the artist whose style is being recognizable and mimicked by the Generative AI art model). To explore the relationship between the Likert scale answers to RQ4_B and the secondary variables, we ran an ordinal logistic regression (OLR), as shown in Figure 7 with the odds ratio of the OLR. Table 4 shows that there is strong evidence of an association (p-value ¡ 0.05) between agreeing that the artist whose style is recognizable in the output should be the owner of the art and the participant’s race and whether they have used AI art models to create art.

### 14 May 2024 v4

@@ 2.4. Problems in the pipeline of Generative AI tools
As discussed earlier, when assessing the impact of an AI model, it extends beyond the usage of the model’s output or its deployment (such as replacing human workers). Every stage involved in developing a Generative AI model can influence the eventual harm caused by the model, thereby affecting its overall impact. Although many stages in this pipeline are typically inaccessible, except for the input and output phases, they still play a significant role. In this section, we outline the pipeline of creating a Generative AI model and focus on one prevalent issue, bias (Zhao et al., 2023), to illustrate how various stages within the pipeline can contribute to the model’s impact.
When developing an AI model, several steps are involved. Firstly, training materials need to be collected, often comprising a subset of real-world artifacts. These materials then undergo an assessment to address issues like poor quality, irrelevance, or over-representation of certain artifacts  (Luccioni and Viviano, 2021; Zhao et al., 2023). Next, the model’s architecture is assembled, followed by the training process. For complex Generative AI models with unpredictable behavior, there is an additional step called alignment. This step aims to constrain the model’s range of potential behaviors to align with the creators’ desired outcomes (Zhao et al., 2023).
@@
Early in the pipeline, the selection of training material significantly shapes the eventual behavior of the model. For instance, many important datasets used in machine learning benchmarks predominantly feature lighter-skinned subjects (Buolamwini and Gebru, 2018), which indirectly introduces bias into the model. This type of global bias cannot be detected by locally examining each piece of training data for harmful content. Furthermore, datasets often contain objectionable and biased content, including racist hate speech and sexually inappropriate images (Luccioni and Viviano, 2021), even after undergoing cleaning procedures.
In the middle of the pipeline, within the internal workings of the trained model, even fundamental and abstract elements like the meanings of common English words have been discovered to encode gender biases (Bolukbasi et al., 2016). As we progress toward the end of the pipeline, the output generated by the model can exhibit biases as well. For example, the Stable Diffusion model disproportionately generates images of white male executives, doctors, lawyers, judges, and law enforcement, and also disproportionately generates images of black and brown criminals and service workers (Ghosh and Caliskan, 2023). Additionally, users have reported that Generative AI tools often generate images where they appear younger, more sexualized, or with lighter skin tones, even when these attributes were not requested or suggested in the input prompt  (Ghosh and Caliskan, 2023). These examples are not unique; similar biases are found across current AI models and products – including those of Generative AI – at every level, from the training data they learned from to their internal representations to their outputs (Luccioni and Viviano, 2021; Ghosh and Caliskan, 2023; Cheng et al., 2023; Zhao et al., 2023).
@@
As a final problem in the pipeline, it is worth noting that these AI tools cannot be created ex nihilo: they are integrated into complex human systems that bring their own forces to bear on the model’s potential impact. There are discrepancies in the demographics of who is wielding the power of these new technologies: who is making them (Stanford Institute for Human-Centered Artificial Intelligence, 2023), profiting off them, using them. This can have many knock-on effects, sometimes through mechanisms difficult to foresee (Even-Tov et al., 2023), with respect to the design, use, and role of these tools in society (Jiang et al., 2023; Stanford Institute for Human-Centered Artificial Intelligence, 2023). In particular, the companies behind these products often employ a great deal of secrecy as to their training data, architecture, or implementation (or all three) (Ho, 2024). As a result, users must interact with the product as a black box (Jiang et al., 2023; Marcus and Southen, 2024); concerned users can only conjecture as to what kind of training data could produce the observed results based on what is known of the basic architecture of the model, which in turn allows argument over the obscure details. The lack of transparency benefits the companies offering these products (with respect to establishing the impact, specifically the harms, of the model) in giving them a way to sidestep moral and ethical concerns (Marcus and Southen, 2024).
3. Methods
3.1. Survey Methodology
Our anonymous original survey of artists is an observational study (N=516) of artists over the age of 18 years that was collected using a convenience sample in phase 1 and supplemented with a Qualtrics survey panel in phase 2 (demographic breakdown can be seen in Table 1. Qualtrics is a reputable academic survey panel platform that has been shown to yield reliable results in comparison to other survey platforms (Boas et al., 2018). For comparison of the distributions of the US art workforce) based on the demographics is presented in Table 4. Incomplete surveys were dropped from the analysis, leaving a final N of 459. Participants provided consent by agreeing to the survey information sheet. They were then prompted to respond to inquiries regarding their artistic identity and financial interactions within the art domain. Participants were presented with a brief introduction to AI-generated art, including an illustrative example, and were asked to indicate their familiarity with AI art tools and models. Subsequently, participants were queried on their perceptions regarding ownership of art produced by AI systems using a 5-point Likert scale. Further questions focused on artists’ preferred compensation criteria for their work being used in AI-generated art models. Participants were also asked to evaluate the impact of AI art models on their field, their perception of these models as potential threats, and whether they believed creators of such models should disclose their data sources using a 5-point Likert scale. Finally, demographic information about the participants’ backgrounds was collected. An appendix that enumerates all survey questions is available in Appendix I. This survey was approved as “exempt” by the University of Vermont Institutional Review Board (IRB).
Phase 2 Sampling: Phase 2 (n = 264) was collected via a targeted survey panel conducted by Qualtrics from September 12-23, 2023 on the Qualtrics online survey platform.
3.2. Data
Our original survey of artists is a convenience survey (N=459).We transformed all survey response variables of interest into numerical form to analyze our survey results (survey questions 8,9,10,13,14,15 as seen in Appendix I). All Likert survey questions were converted from ‘Strongly Agree,’ ‘Agree,’ ‘Neutral,’ ‘Disagree,’ and ‘Strongly Disagree’ to an ordinal scale of 1,2,3,4,5. For our ordinal logistic regression analysis, we combine ‘Strongly Agree’ and ‘Agree’ to be equal to 3, ’Neutral’ to be equal to 2, and ‘Disagree’ and ‘Strongly Disagree’ to be equal to 1. In the original logistic regression, we reduced the dimensionality of the Likert scale to investigate the positive, neutral, and negative sentiment of the response questions in relation to our secondary variables.
All nominal and categorical variables (survey questions 1,2,3,4,5,11,12,16,17,18,19, as seen in Appendix I) were transformed into binary variables. Survey questions that included an “other” option or write-in answers were also categorized (e.g., in question 2, we added an additional art practice, which included digital art from the write-in answers).
Demographic questions (survey questions 16, 17, 18, 19 as seen in Appendix I) included asking survey participants their gender identity to identify whether they were white or a person of color, choosing their age range and country of residence. All demographic questions were transformed into binary variables.
3.3. Analytical Methods
Odds ratio In general, the odds of an event are the probability of success over the probability of failure. This can be written as the following equation:
(1) $\frac{p}{1-p}$
Ordinal Logistic Regression Ordinal logistic regression (OLR) is a generalization of regression (either multiple linear or binomial logistic) used for an ordered categorical response variable and any number of categorical or numeric secondary variables. OLR finds the logit (also known as log-odds) of the dependent variable (Y) being less than or equal to a certain level (denoted j) (Bruin, 2011b). which can be written as:
(2) $logit(P(Y\leq j))=log\frac{P(Y\leq j)}{P(Y>j)}$
It is important to note that OLR assumes proportional odds, meaning the effect of the secondary variable(s) is the same between all levels of the dependent variable.
@@ 4. Results
Table 1. Descriptive statistics for our survey participants, N=459. A detailed table of descriptive statistics can be seen in the Appendix in Table 5
@@
459 participants were anonymously surveyed, yielding a sample reflective of the demographics (gender, ethnicity, age) of the broader American adult art workforce (see Table 1). We conducted this observational study on the opinions of artists to explore the dynamic interplay between their perspectives on the potential benefits and risks of Generative AI art. This investigation delves into artists’ viewpoints regarding the utility and risks associated with Generative AI art models, ethical considerations in disclosing artistic works used in AI art training models, the ownership and rights pertaining to AI art derivatives, and equitable compensation. Our results indicate a strong sentiment among artists in favor of disclosing what art is being used in AI training (RQ3), advocating against generative AI companies retaining ownership of AI-generated outputs (RQ4), and expressing concerns regarding AI’s impact on the art industry’s workforce (RQ1-2) and who profits from their art.
In the analysis of survey data pertaining to our primary research inquiries (see Table 2), 80.17% of the survey participants agree that model creators should be required to disclose in detail what art and images they use to train their AI models. We also find that 61.87% of survey participants agree that AI models are a threat to art workers. In contrast, 44.88% of survey participants agree that AI art models are a positive development in the field of art, notably with 27.02% of artists selecting neutral on the Likert scale for this question. Interestingly, we see an overlap between artists who agree that AI art is a positive development and those who also agree that it is a threat to art workers; 22.03% of survey participants agreed that both AI was a positive development and a threat, which presents an interesting tension which highlights that the sentiments of artists are nuanced and some hold complex opinions towards generative AI.
We examine a survey question prompting participants to consider a scenario where an AI art model was employed by another individual to create artwork distinctly resembling the participant’s style and reflect on the rightful ownership of the resulting work and any derivative creations. Only 26.80% of the participants agreed that the work and its derivatives be considered the property of the AI model creators, while 39.22% of participants agreed that the work and its derivatives be considered the property of the person who used the AI model to generate the artwork. Finally, 41.39% of the participants agreed that the work and its derivatives be considered the property of the artists whose style is being represented by the AI model output. This result presents an interesting tension as well between artists who think the ownership should belong to the original inspiration of the AI art while others may also want to retain the ownership of the resulting work. We see in our results that artists who have used AI art models are more likely to agree that the resulting work should be considered the property of the person who used the model to create the work.
In the following sections, we will present results that identify the variables most strongly associated with survey participants who expressed agreement with the primary research questions.
@@
Figure 2. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ1 (Generative AI art models as a threat to art workers) and other survey variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value $<$ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
@@
Figure 3. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ2 (Generative AI art models as a positive development) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value $<$ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
@@
Figure 4. The ordinal logistic regression results show the association between answering ‘Agree’ or ‘Strongly Agree’ to RQ3 (Should model creators be required to disclose in detail what art & images they use to train their AI models) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value $<$ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.
RQ1: Do artists see Generative AI models as a threat to art workers?
This research question investigates artists’ perceptions of Generative AI models as potential threats to art workers. Our findings reveal that 61.87% of survey participants hold the view that Generative AI models pose a threat to art workers. Thus, while a majority of artists perceive AI as a threat, it is not an overwhelming majority. This finding is related to survey question no. 15 (see Appendix I for survey questions).
An ordinal logistic regression (OLR) analysis was performed to examine the relationship between the Likert scale responses to RQ1 (dependent variable) and the secondary variables (all odds ratios can be seen in Appendix III), focusing on the odds ratios derived from the OLR model (as seen in Figure 2).
The results demonstrate a significant relationship (p-value $<$ 0.05, all detailed odds ratios can be seen in Appendix III) between agreeing (represents a combination of both the responses ‘Agree’ and ‘Strongly Agree’ in the Likert scale) that Generative AI models are a threat to art workers and a subset of our secondary variables. Here, we see that artists ages 18-29 are more likely to agree that there is a threat and that artists who have used Generative AI art models to create art are less likely to think that it is a threat to the art workforce.
RQ2: Do artists see Generative AI art models as a positive development in the field of art?
This research question investigates artists’ perceptions of AI models as positive development to the field of art. We ran the same analysis for our second research question and found that 44.88% of the survey participants agree that Generative AI art models are a positive development in the field of art. Notably, 27.02% of artists selected ’neutral’ on the Likert scale for this question. This analysis pertains to survey question number 14 (see Appendix I for survey questions).
We conducted an ordinal logistic regression (OLR) analysis to investigate the association between Likert scale responses to RQ2 (dependent variable) and the secondary variables with a focus on the odds ratios obtained from the OLR model (as seen in Figure 3). Our analysis indicates a strong relationship (p-value $<$ 0.05, all OLR detailed results can be seen in Appendix III) between agreeing that Generative AI art models are a positive development in the field of art and a subset of our secondary variables. Here, we see that 30-49 years olds, males, persons of color, participants who have used AI art models, and participants who said they would donate their artwork to train models were more likely to agree that generative AI models are a positive development for the field of art. Whereas makers (note that makers were a small n (n=5) with a fairly large range in the confidence interval bars represented in Figure 3), females, and those who identify as artists were less likely to agree that AI models are a positive development for the field of art.
RQ3: What types of disclosure for the use of artwork to train an AI art model do artists consider fair?
This research question investigates artists’ perceptions of what types of disclosure are fair when using their artwork to train an AI art model. Our results show a consensus among a large majority of artists that transparency and disclosure regarding the specific images used in training an AI model are essential prerequisites for them to deem the model fair and acceptable. We see from our survey results that 80.17% of the survey participants agree that creators of AI models should disclose the art and images used to train their models in detail. The analysis for RQ3 is related to survey question 13 (all survey questions can be seen in Appendix I), which we used to run an ordinal logistic regression (OLR) to explore the relationship between the Likert scale answers to RQ3 and our secondary variables. The results of the OLR, displayed in Figure 4 as the odds ratio, show a strong relationship (p-value $<$ 0.05, all odds ratios can be seen in Appendix III) between agreeing with the disclosure requirement and a subset of our secondary variables.
Here, we see that participants with a general familiarity with AI art or were between the ages of 18-49 were more likely to agree that AI model creators should be required to disclose their training data in detail. Participants who expressed willingness to accept compensation in the form of a share of the derivatives or profits generated from AI artwork inspired by their work, along with a portion of the model creators’ profits, also agreed that disclosure should be required. Participants who identified as illustrators (note that illustrators were a small n (n=9) with a fairly large range in the confidence interval bars represented in Figure 4) or were between the ages of 50-64 years old were less likely to agree that disclosure should be required.
@@
Figure 5. The ordinal logistic regression results show the association between the answering ‘Agree’ or ‘Strongly Agree’ to RQ4A (If an AI art model was used by someone else to produce artwork recognizably in your style (e.g., in the style of Van Gogh) Should that work and its derivatives be considered the property of the person who used the AI model to generate the artwork?) and survey secondary variables. Results are represented by odds ratios on a log scale on the x-axis, with bars conveying a 95% confidence interval and red indicating a p-value $<$ 0.05 and a confidence interval that does not cross the threshold of the dotted line in the figure.